uestion 6 (1 point) In its 2018 fiscal year, the data storage company, NetApp Inc.,, reported that it had 267.9 million shares of common stock outstanding, trading at a price of about $68 per share. On the firms balance sheet, the value of common stock equity was reported as $2.067 billion. NetApps market/book ratio was A) 8.8 B) 7.7 C) 68 D) greater than 100
Answer #1

Market value = Share price * Shares outstanding

Market Value = 68 * 267,900,000 = 18,217,200,000

Book Value = 2,067,000,000

Market to book ratio = 18,217,200,000 / 2,067,000,000 = 8.8